OpenStudy (anonymous):

Can someone explain to me what closure property is and how to identify it from an equations?

Directrix (directrix):

Do you have a specific problem in mind? If so, please post it.

OpenStudy (anonymous):

Not really. I have a math teacher that assumed we learned this already and it's not in the math textbook. I just need help with the problems where you identify the property ex. 3+(4+5)=3+(5+4) Commutative property

Directrix (directrix):

Start with the properties attached in the file.

Directrix (directrix):

After you read them, go to the practice site at the following link: http://www.regentsprep.org/Regents/math/ALGEBRA/AN1/propPrac.htm

OpenStudy (anonymous):

I know these properties already. The only property I don't understand is the closure property

Directrix (directrix):

Okay, I want to ask you to answer this question first. Thanks. Name the property: For real numbers, a, b, and c, give the property that justifies the following: a + (b + c) = a + (c + b)

OpenStudy (anonymous):

Commutative Property because the values contained in the parenthesis stay the same, right?

Directrix (directrix):

Yes, very good. Most people focus on the three elements and answer "associative." Now for closure. See the two attached files.

OpenStudy (anonymous):

Done. How would the problems look like?

Directrix (directrix):

Question 1: Is the set of negative integers closed under the operation of multiplication? Questions 2: Is the finite set { -1, 0, 1 } closed under the operation of division?

OpenStudy (anonymous):

Q1: No Q2: No

OpenStudy (anonymous):

But are there problems that look like For all real numbers a and b, {x|x.....? My teacher seems to love those.

Directrix (directrix):

Your answers to Q1 and Q2 are correct. That is set notation your teacher is using. If you would look in your notes and find one of those problems and post it here, that would be good. We could discuss it.

OpenStudy (anonymous):

I'm not sure if I copied it down, but I'll check >.<

OpenStudy (anonymous):

I can't find it, but the problems after the closure one all began with something like